# Enhancing Agentic textual Graph Retrieval with Synthetic Stepwise Supervision

### v1

@@ 4.1 Experimental Setup
Datasets. We evaluate Graph-S3 on three widely used textual graph QA benchmarks. WebQSP (Yih et al., 2015) consists of real-world questions annotated with SPARQL queries against Freebase, primarily involving one- or two-hop reasoning. CWQ (Talmor & Berant, 2018) extends WebQSP with more complex multi-hop questions, posing a greater challenge for long reasoning chains. MetaQA (Zhang et al., 2018) is a movie-domain benchmark containing 135k triples and 43k entities, designed to evaluate multi-hop reasoning in a closed domain. Following prior work (Chen et al., 2024b), we report accuracy (Acc) and $\mathbf{F_{1}}$ score as evaluation metrics.
Baselines. To validate the effectiveness of our approach, we compare with several representative graph retrieval methods. We additionally evaluate the model’s inherent graph understanding capability through two configurations: (1) the no graph setting, where the model processes the query without any graph input, and (2) the no retriever setting, where the model receives the entire graph structure directly as input. For traditional RAG, we implemented a multi-hop method where the model retrieves the most relevant graph nodes for the current query and then performs a k-hop expansion to collect information for answer generation. We further compare with representative graph retrievers, including Think-on-Graph (ToG) (Sun et al., 2023), LightRAG (Guo et al., 2024b), G-Retriever (He et al., 2024) and agentic graph retrieval system KG-Agent (Jiang et al., 2024).
Implementation Details. In our experiments, we primarily employed the Llama3.1-8B (Dubey et al., 2024) and Qwen3-8B (Yang et al., 2025) models. Our data synthesis pipeline produced 9,035 SFT and 3,504 RL training instances; with this data, the Qwen3-8B model was trained for 3 SFT and 15 RL epochs on 8 A100 GPUs, requiring 32 hours in total. Further training details are provided in the Appendix A.3.
@@ 4.2 Main Results
The results of WebQSP, CWQ, and MetaQA are summarized in Table 1. In general, our framework achieves the best performance among all compared methods, demonstrating the effectiveness of combining two-stage training with interactive retrieval. Compared with the no-retriever configuration, where the entire graph is directly fed into the LLM, retrieval-based methods consistently achieve higher accuracy, confirming that selective subgraph retrieval is essential since full graphs exceed the effective processing capacity of LLMs. Relative to $k$-hop expansion approaches, Graph-S3 yields clear improvements, particularly on multi-hop benchmarks, showing that interactive retrieval can effectively filter relevant relations while avoiding redundant context. Furthermore, against training-free baselines such as ToG and LightRAG, our model delivers substantial gains, highlighting the importance of stepwise synthetic supervision and reinforcement learning in enhancing the reasoning ability of the retriever. Compared with trained retrievers such as G-Retriever and existing agentic graph retrieval frameworks like KG-Agent, our approach further improves multi-hop reasoning performance, indicating that the introduction of reinforcement learning and stepwise supervision enables the retriever to acquire stronger reasoning ability beyond representation learning.
@@ 4.3 In-depth Analysis
Graph-S3 44.29 58.45 23.62 30.44 82.77 92.04 63.17 76.18 14.70 36.29
Graph-S3 w/o SFT 31.64 44.41 7.74 8.77 81.27 89.38 46.30 54.12 2.07 4.98
Graph-S3 w/o RL 41.77 53.02 13.39 15.97 71.97 80.09 35.93 45.25 5.73 11.46
Graph-S3 w/o interactive 28.87 41.48 19.85 26.01 59.50 69.54 1.83 7.44 3.30 18.61
Graph-S3 w/o trajectory refinement 16.46 19.24 4.12 4.87 39.47 41.06 4.01 6.10 1.34 1.80
Table 2: Results of abalation studies.
@@ 4.3.1 Ablation Study
Graph-S3 w/o step supervision 41.83 53.87 13.47 16.40 72.63 81.12 34.97 45.14 6.43 11.34
Graph-S3 44.29 58.45 23.62 30.44 82.77 92.04 63.17 76.18 14.70 36.29
@@ 4.3.2 Effectiveness of Stepwise Supervision
Figure 3: Number of retrieved graph triples in Graph-S3 and baselines on correct answers.
@@ 4.3.3 Effective Information Quantification Analysis
To evaluate the efficiency of Graph-S3 in retrieving concise yet effective information, we compare it with baseline approaches by measuring the number of triples required to produce correct answers (see Figure 3). Unlike traditional methods that often retrieve large amounts of redundant information, our approach significantly reduces retrieval size while maintaining higher accuracy. In particular, our experiments show that Graph-S3 requires only 11.44% of the triples retrieved by G-Retriever on average, yet still achieves superior accuracy. These results highlight the framework’s ability to balance search depth with precision, thereby reducing redundancy.
@@ 5 Conclusion
We investigated the limitations of existing retrieval-augmented generation methods on textual graphs, highlighting their reliance on outcome-based supervision and their tendency to produce redundant or incomplete subgraphs. To overcome these challenges, we proposed a framework that integrates three key innovations: (1) A pipeline for high-quality, stepwise-supervised data synthesis; (2) Two-stage training (SFT then RL) with process-level rewards; (3) Fine-grained, interactive retrieval over textual graphs. Extensive experiments on WebQSP, CWQ, and MetaQA demonstrate that our approach consistently improves both accuracy and $F_{1}$, validating the effectiveness of synthetic stepwise supervision and the proposed training strategy for enhancing interactive graph retrieval.
@@ 6 Acknowledgments
This work is supported by National Natural Science Foundation of China (Grant No.62272261), Tsinghua University (AIR)–AsiaInfo Technologies (China) Inc. Joint Research Center, and Wuxi Research Institute of Applied Technologies, Tsinghua University.
Appendix A Appendix
A.1 Training Details
Implementation Details. For data generation, we apply our proposed data synthesis pipeline, producing a total of 9,035 training instances for SFT and 3,504 instances for RL. In the SFT stage, we fine-tune the Qwen3-8B with a learning rate of $1\times 10^{-4}$ for $3$ epochs. In the RL stage, we adopt the GRPO algorithm with a batch size of $512$, $15$ training epochs, a learning rate of $1\times 10^{-5}$, a value clipping range of $0.5$, and a KL divergence coefficient of $0.001$. The entire RL training phase takes approximately $32$ hours on $8$ NVIDIA A100 80GB GPUs.
@@ A.1 Training Details
A.2 Prompt of Interactive Graph Retrieval
@@
Prompts for Interactive Graph Retriever You are an intelligent agent skilled in exploring Knowledge Graphs, with strong reasoning abilities. Your task is to perform question answering over a Knowledge Graph by gradually exploring it. You should start from the entities mentioned in the question and explore the graph step by step until you gather enough information to answer the question. Your task follows these steps: 1. Understand the Question 2. Analyze the Action History and Current Graph State 3. Choose the Next Action** from the following options: "Explore Entity": Explore all triples directly connected to a given entity in the Knowledge Graph. "Choose Relation": Select the triple(s) from the explored information that are most relevant to the question. Attention: Only the triples included in the "Objects" field of the "Choose Relation" step will be retained in the future "Current Graph State". So You must filter and retain the information useful for answering the question or for further exploration. "Finish": Choose this action when you believe you have gathered sufficient information to answer the question. Your final answers should be included in the "Objects" field. 4. Select the Objects: Depending on the action, provide the relevant entity or triple(s). Attention: All objects must come from the "Entities in Question" or the current "Current Graph State". Do not create new entities or relations. If the action is "Explore Entity", "Objects": ["EntityA", "EntityB"] If the action is "Choose Relation", "Objects": ["(Subject1, Relation1, Object1)", "(Subject2, Relation2, Object2)"] If the action is "Finish", "Objects": ["Answer1", "Answer2"] 5. Output your response in JSON format, and include a **detailed thought process explaining your reasoning at this step. --- Question: Entities in Question: Current Graph State: Action History: --- Please respond using the following format: Thought Process: <Provide a step-by-step analysis> Action Decision: ‘‘‘json "Action": "<The type of action you are taking: ’Explore Entity’ | ’Choose Relation’ | ’Finish’>", "Objects": [<The entities or triples>]
A.3 An Example of The Reasoning Process of Graph-S3
@@
"step 5, Explore Entity, Objects: [’Ugetsu, Sansho the Bailiff, The Crucified Lovers, The 47 Ronin, Osaka Elegy, Sisters of the Gion, Utamaro and His Five Women’]"
@@
"step 5, Explore Entity, Objects: [’Ugetsu, Sansho the Bailiff, The Crucified Lovers, The 47 Ronin, Osaka Elegy, Sisters of the Gion, Utamaro and His Five Women’]",
@@
A.4 The Use of Large Language Models
In this work, large language models (LLMs) were utilized solely to assist in polishing the academic writing. Specifically, the LLM (e.g., [DeepSeek]) was used to improve the fluency and clarity of the textual descriptions in Section 3 (Method) and Section 4 (Experiments). All technical content, including the core ideas, algorithms, experimental results, and analyses, was conceived and developed solely by the authors. The LLM acted strictly as a writing tool and did not contribute to the scientific reasoning or intellectual substance of the research.

### v3

@@ 4.1 Experimental Setup
Datasets. We evaluate Graph-$S^{3}$ on three widely used textual graph QA benchmarks. WebQSP (Yih et al., 2015) consists of real-world questions annotated with SPARQL queries against Freebase, primarily involving one- or two-hop reasoning. CWQ (Talmor and Berant, 2018) extends WebQSP with more complex multi-hop questions, posing a greater challenge for long reasoning chains. MetaQA (Zhang et al., 2018) is a movie-domain benchmark containing 135k triples and 43k entities, designed to evaluate multi-hop reasoning in a closed domain. Following prior work (Chen et al., 2024), we report accuracy (Acc) and $\mathbf{F_{1}}$ score as evaluation metrics.
Baselines. To validate the effectiveness of our approach, we compare with several representative graph retrieval methods. We additionally evaluate the model’s inherent graph understanding capability through two configurations: (1) the no graph setting, where the model processes the query without any graph input, and (2) the no retriever setting, where the model receives the entire graph structure directly as input.
For traditional RAG, we implemented a multi-hop method where the model retrieves the most relevant graph nodes for the current query and then performs a k-hop expansion to collect information for answer generation. We further compare with representative graph retrievers, including Think-on-Graph (ToG) (Sun et al., 2023), LightRAG (Guo et al., 2024b), G-Retriever (He et al., 2024).
Implementation Details are provided in the Appendix A.
@@ 4.2 Main Results
The results of WebQSP, CWQ, and MetaQA are summarized in Table 1. In general, our framework achieves the best performance among all compared methods, demonstrating the effectiveness of combining two-stage training with interactive retrieval1. We break down the analysis into several key observations.
Superiority in Complex Multi-hop Reasoning. Graph-$S^{3}$ exhibits a widening performance gap over baselines as reasoning complexity increases. On the MetaQA benchmark, while competitive on 1-hop questions, our method dominates on multi-hop tasks. Specifically, on MetaQA 3-hop, Graph-$S^{3}$ significantly outperforming the next best agentic baseline and far surpassing traditional $k$-hop RAG methods which struggle to maintain coherence over long chains.
Effectiveness Against Training-Free Baselines. Against sophisticated training-free baselines like ToG and LightRAG, our model delivers substantial gains across all datasets. This highlights that while LLM-based planning (ToG) or structural awareness (LightRAG) is beneficial, it is insufficient without the dedicated fine-tuning and reinforcement learning provided by our pipeline to align the retriever with graph-specific reasoning patterns.
Advantage Over Trained Retrievers. Compared to supervised retrievers like G-Retriever, Graph-$S^{3}$ consistently achieves higher $F_{1}$ scores, indicating better precision in answer generation. This improvement suggests that our interactive pruning mechanism effectively balances recall and precision—retrieving enough context to answer correctly without overwhelming the generator with noise, a common pitfall for G-Retriever which retrieves larger, noisier subgraphs.
Robustness Across Backbones. The performance gains are consistent across different LLM backbones and fine-tuning strategies. Notably, even when using a fine-tuned backbone, Graph-$S^{3}$ maintains its lead, achieving the highest scores on all benchmarks. This demonstrates that our method’s benefits are orthogonal to the underlying model’s capability and can serve as a universal enhancement for graph-based QA systems.
To gain deeper insights into the quantitative improvements, we analyzed specific reasoning examples to verify the effectiveness of our interactive pruning and stepwise supervision. A detailed presentation of these case studies and the corresponding reasoning processes of Graph-$S^{3}$ can be found in Appendix C.
@@ 4.3 In-depth Analysis
Graph-$S^{3}$ 44.29 58.45 23.62 30.73 82.77 92.04 63.17 76.18 14.70 36.29
w/o SFT 31.64 44.41 7.74 8.77 81.27 89.38 46.30 54.12 2.07 4.98
w/o RL 41.77 53.02 13.39 15.97 71.97 80.09 35.93 45.25 5.73 11.46
w/o interactive 28.87 41.48 19.85 26.01 59.50 69.54 1.83 7.44 3.30 18.61
w/o traj. refine 16.46 19.24 4.12 4.87 39.47 41.06 4.01 6.10 1.34 1.80
Table 2: Results of ablation studies.
@@ 4.3.1 Ablation Study
w/o step supervision 41.83 53.87 13.47 16.40 72.63 81.12 34.97 45.14 6.43 11.34
Graph-$S^{3}$ 44.29 58.45 23.62 30.73 82.77 92.04 63.17 76.18 14.70 36.29
@@ 4.3.2 Effectiveness of Stepwise Supervision
Figure 3: Number of retrieved graph triples in Graph-$S^{3}$ and baselines on correct answers.
@@ 4.3.3 Effective Information Quantification Analysis
To evaluate the efficiency of Graph-$S^{3}$ in retrieving concise yet effective information, we compare it with baseline approaches by measuring the number of triples required to produce correct answers (see Figure 3). Unlike traditional methods that often retrieve large amounts of redundant information, our approach significantly reduces retrieval size while maintaining higher accuracy. In particular, our experiments show that Graph-$S^{3}$ requires only 11.44% of the triples retrieved by G-Retriever on average, yet still achieves superior accuracy. These results highlight the framework’s ability to balance search depth with precision, thereby reducing redundancy.
@@ 5 Conclusion
We investigated the limitations of existing retrieval-augmented generation methods on textual graphs, highlighting their reliance on outcome-based supervision and their tendency to produce redundant or incomplete subgraphs. To overcome these challenges, we proposed a framework that integrates three key innovations:
(1) A pipeline for high-quality, stepwise-supervised data synthesis; (2) Two-stage training (SFT then RL) with process-level rewards; (3) Fine-grained, interactive retrieval over textual graphs. Extensive experiments on WebQSP, CWQ, and MetaQA demonstrate that our approach consistently improves both accuracy and $F_{1}$, validating the effectiveness of synthetic stepwise supervision and the proposed training strategy for enhancing interactive graph retrieval. These findings underscore the potential of fine-grained process supervision in bridging the gap between LLMs and structured knowledge, paving the way for more reliable and interpretable agentic reasoning systems.
6 Limitation
While this work has focused on validating Graph-S³ on standard textual graph benchmarks, future research could explore extending the interactive retrieval framework to diverse problem scenarios. Additionally, with the rapid advancement of agent technology, the current framework can be further enhanced by designing a broader spectrum of actions and integrating external tools to tackle increasingly complex reasoning tasks.
@@ 7 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No. 62272261, Wuxi Research Institute of Applied Technologies, Tsinghua University under Grant 20242001120, Tsinghua University (AIR)–AsiaInfo Technologies (China) Inc. Joint Research Center and a grant from GDS Holdings Limited.
Appendix A Training Details
Implementation Details. In our experiments, we primarily employed the Llama3.1-8B (Dubey et al., 2024) and Qwen3-8B (Yang et al., 2025) models. Our data synthesis pipeline produced 9,035 SFT and 3,504 RL training instances; with this data, the Qwen3-8B model was trained for 3 SFT and 15 RL epochs on 8 A100 GPUs, requiring 32 hours in total.
For data generation, we apply our proposed data synthesis pipeline, producing a total of 9,035 training instances for SFT and 3,504 instances for RL. In the SFT stage, we fine-tune the Qwen3-8B with a learning rate of $1\times 10^{-4}$ for $3$ epochs. In the RL stage, we adopt the GRPO algorithm with a batch size of $512$, $15$ training epochs, a learning rate of $1\times 10^{-5}$, a value clipping range of $0.5$, and a KL divergence coefficient of $0.001$. The entire RL training phase takes approximately $32$ hours on $8$ NVIDIA A100 80GB GPUs.
@@ Appendix A Training Details
Appendix B Prompt of Interactive Graph Retrieval
@@
Prompts for Interactive Graph Retriever You are an intelligent agent skilled in exploring Knowledge Graphs, with strong reasoning abilities. Your task is to perform question answering over a Knowledge Graph by gradually exploring it. You should start from the entities mentioned in the question and explore the graph step by step until you gather enough information to answer the question. Your task follows these steps: 1. Understand the Question 2. Analyze the Action History and Current Graph State 3. Choose the Next Action from the following options: "Explore Entity": Explore all triples directly connected to a given entity in the Knowledge Graph. "Choose Relation": Select the triple(s) from the explored information that are most relevant to the question.
Attention: Only the triples included in the "Objects" field of the "Choose Relation" step will be retained in the future "Current Graph State". So You must filter and retain the information useful for answering the question or for further exploration.
"Finish": Choose this action when you believe you have gathered sufficient information to answer the question. Your final answers should be included in the "Objects" field.
4. Select the Objects: Depending on the action, provide the relevant entity or triple(s).
Attention: All objects must come from the "Entities in Question" or the current "Current Graph State". Do not create new entities or relations.
Examples: If "Explore Entity": "Objects": ["EntityA", "EntityB"] If "Choose Relation": "Objects": ["(Sub1, Rel1, Obj1)", ...] If "Finish": "Objects": ["Answer1", "Answer2"]
5. Output your response in JSON format, and include a detailed thought process explaining your reasoning at this step.
  Question:
Entities in Question:
Current Graph State:
Action History:
  Please respond using the following format: Thought Process:
<Provide a step-by-step analysis>
Action Decision:
‘‘‘json
"Action": "<The type of action: ’Explore Entity’ | ’Choose Relation’ | ’Finish’>",
"Objects": [<The entities or triples>]
Appendix C Case Study
Case 1 Question: When were the movies directed by [Sarah Polley] released? True Answer: 2011, 2006, 2012 Action History: Step 1, Explore Entity, Objects: [’Sarah Polley’] Step 2, Choose Relation, Objects: [’(Away from Her, directed_by, Sarah Polley)’, ’(Stories We Tell, directed_by, Sarah Polley)’, ’(Take This Waltz, directed_by, Sarah Polley)’] Step 3, Explore Entity, Objects: [’Stories We Tell’, ’Away from Her’, ’Take This Waltz’] Step 4, Choose Relation, Objects: [’(Away from Her, release_year, 2006)’, ’(Stories We Tell, release_year, 2012)’] Step 5, Explore Entity, Objects: [’Take This Waltz’] Step 6, Choose Relation, Objects: [’(Take This Waltz, release_year, 2011)’] Step 7, Finish, Answer: [’2006’, ’2011’, ’2012’] Retrieved Triples: (Away from Her, release_year, 2006), (Away from Her, directed_by, Sarah Polley), (Take This Waltz, release_year, 2011), (Stories We Tell, directed_by, Sarah Polley), (Stories We Tell, release_year, 2012), (Take This Waltz, directed_by, Sarah Polley)
To demonstrate the efficacy of our interactive pruning mechanism against high-degree entities, we analyze the query: "When were the movies directed by [Sarah Polley] released?" The entity "Sarah Polley" presents a significant challenge as a "hub node" with dual roles: she is a prolific actress with numerous connections but has directed comparatively fewer films.
Graph-$S^{3}$ Success via Structure-Aware Pruning: As shown in the trajectory, Graph-$S^{3}$ effectively overcomes the noise of Sarah Polley’s acting career. In Step 2, the agent leverages the "Choose Relation" action to strictly select triples containing the "directed_by" relation, while deliberately pruning the dense subgraph of movies where she appears merely as an actress. This aligns with our framework’s design to "avoid introducing redundant context into the LLM" and demonstrates the value of stepwise supervision in learning to distinguish task-relevant paths from high-degree noise. Consequently, the agent precisely navigates to the correct release years.
In contrast, baseline methods exhibited failures characteristic of shallow embedding similarity. The standard RAG suffered from severe lexical confusion. It retrieved unrelated entities such as "Pollyanna" and "Polly of the Circus", erroneously matching the token "Polly" with "Polley." This confirms that similarity-based matching often yields noisy results by ignoring precise relational constraints. G-Retriever failed to filter the entity’s large neighborhood, retrieving generic attributes instead of the specific directorial path. This illustrates the limitation of one-shot retrievers that retrieve large neighborhoods discarding the graph’s relational structure, whereas Graph-$S^{3}$’s interactive policy successfully identified the sparse signal within a noisy subgraph.
Case 2 Question: The screenwriter [Jane Anderson] co-wrote movies with whom? True Answer: Anne Heche, Terry Ryan Action History: Step 1, Explore Entity, Objects: [’Jane Anderson’] Step 2, Choose Relation, Objects: [’(If These Walls Could Talk 2, written_by, Jane Anderson)’, ’(Normal, written_by, Jane Anderson)’, ’(The Prize Winner of Defiance, Ohio, written_by, Jane Anderson)’] Step 3, Explore Entity, Objects: [’If These Walls Could Talk 2’, ’Normal’, ’The Prize Winner of Defiance, Ohio’] Step 4, Choose Relation, Objects: [’(If These Walls Could Talk 2, written_by, Jane Anderson)’, ’(Normal, written_by, Jane Anderson)’, ’(The Prize Winner of Defiance, Ohio, written_by, Jane Anderson)’] Step 5, Finish, Answer: [’If These Walls Could Talk 2’, ’Normal’, ’The Prize Winner of Defiance, Ohio’] Retrieved Triples: (The Prize Winner of Defiance, Ohio, written_by, Jane Anderson), (Normal, written_by, Jane Anderson), (If These Walls Could Talk 2, written_by, Jane Anderson)
To identify limitations in current reasoning policies, we analyze a failure case involving the collaborative query: "The screenwriter [Jane Anderson] co-wrote movies with whom?" The trajectory of the agent reveals a critical flaw in the processing of "sibling" relationships. In Steps 1 through 3, the agent correctly identified the intermediate nodes, successfully retrieving movies written by Jane Anderson.
The reasoning collapsed at Step 4. When expanding from the movie nodes to identify other writers, the agent exhibited a strong "self-verification bias." Instead of selecting edges pointing to new entities (the co-writers, such as Anne Heche), the agent selected the "written_by" relations pointing back to the query entity, Jane Anderson. Consequently, instead of discovering the target "sibling" nodes, the agent entered a redundant verification loop. It merely reconfirmed that Jane Anderson wrote the movies and prematurely terminated the search, outputting the movie titles instead of the co-writers’ names. The model assigns disproportionate attention weight to entities explicitly named in the query, causing it to overlook unmentioned but structurally relevant entities.
Appendix D An Example of The Reasoning Process of Graph-$S^{3}$
@@
"step 5, Explore Entity, Objects: [’Ugetsu’, ’Sansho the Bailiff’, ’The Crucified Lovers’, ’The 47 Ronin’, ’Osaka Elegy’, ’Sisters of the Gion’, ’Utamaro and His Five Women’]"
@@
"step 5, Explore Entity, Objects: [’Ugetsu’, ’Sansho the Bailiff’, ’The Crucified Lovers’, ’The 47 Ronin’, ’Osaka Elegy’, ’Sisters of the Gion’, ’Utamaro and His Five Women’]",